OpenStudy (anonymous):

Lisa Chen invested $24,000, part at 8% and the rest at 7.2%. How much did she invest at each rate if her income from the 8% investment is two thirds that of the 7.2% investment?

OpenStudy (anonymous):

How much of an 18% solution of sulfuric acid should be added to 360 ml of a 10% solution to obtain a 15% solution?

OpenStudy (amistre64):

e+r = 24000 ; e = 24000 - r e(.08) = 2r(.072)/3

OpenStudy (amistre64):

im get r = 15000; but id have to dbl chk that

OpenStudy (anonymous):

$15,000 at 7.2% and $9,000 at 8% thats what i got

OpenStudy (amistre64):

15000 + 9000 doesnt = 45000 tho :)

OpenStudy (amistre64):

...... nvr mind; my mind was lost lol

OpenStudy (amistre64):

15 at 7.2 and 9 at 8% looks good

OpenStudy (anonymous):

thanks bro!